# Cycle-Aware ZZ Crosstalk Mitigation on Quantum Hardware

### v1

@@ 7.2 RQ2 — Qubit Connectivity
Interestingly, our results suggest that CAZZO performs particularly well in environments with Low Connectivity qubit architectures. The low connectivity structures, such as those found in IBMQ-Brisbane and Rigetti’s Aspen-M, tend to benefit more from intelligent gate scheduling, as the physical qubit interactions are constrained by the hardware topology. CAZZO’s ability to optimize scheduling in such cases leads to better parallelism and resource utilization, as demonstrated by the greater speedups observed on these devices. In contrast, Linear Nearest Neighbor architectures [23], which inherently limit gate concurrency, show less dramatic improvements in execution time, though CAZZO still provides notable gains.
@@ 7.3 RQ3 — Fidelity Maintenance
Figure 10 demonstrates CAZZO’s ability to maintain computational fidelity while achieving significant speedups across five benchmark circuits on IBMQ Brisbane. For the bb84_n8 and qec_sm_n5 benchmarks, CAZZO matches ZZXSched’s exact fidelity values. The lpn_n5 circuit reveals CAZZO’s error resilience, maintaining 64% fidelity compared to the Standard scheduler’s 50%, a 14% relative improvement.
Our results disprove the assumption that reduced circuit duration necessarily increases error rates. CAZZO maintains ZZXSched’s error suppression capability (calculated through fidelity ratio analysis) — achieving an optimal balance between speed and accuracy for practical quantum applications.
@@ 7.3.1 Trade-offs Between Density and Fidelity
Our analysis suggests a subtle trade-off between circuit density and fidelity. As CAZZO increases the quantum circuit density by scheduling gates more compactly in time, it reduces the overall execution time, which benefits fidelity to the decreased exposure to decoherence. However, the denser packing of gates also increases the likelihood of crosstalk and other noise sources. The key takeaway from our experiments is that CAZZO strikes a favorable balance, achieving significant reductions in execution time while maintaining high fidelity.
@@ 7.4 RQ4 — Scalability Analysis
Figure 11 presents a scalability comparison of CAZZO versus ZZXSched across four different quantum architectures as the number of qubits increases. In smaller circuits, the gap between CAZZO and ZZXSched is relatively narrow; however, as circuit size grows, CAZZO demonstrates a consistently lower execution time. This trend indicates better scalability, suggesting that CAZZO more effectively mitigates crosstalk and schedules gates in parallel without incurring significant overhead. Although the absolute times vary among architectures due to differences in gate durations and connectivity, CAZZO maintains its advantage in each case.
@@ 9 Conclusions
We formally defined the cycle-aware ZZ crosstalk mitigation problem and proposed a corresponding optimization algorithm. CAZZO balances ZZ crosstalk and parallelism while maximizing time resources in quantum circuits. It uses a new data structure, the Time and Distance Dependency Graph (TDDG), to capture dependencies and adjust gate execution based on varying durations. We assessed CAZZO through simulations on 72 benchmarks from the QASMBench suite and real-device experiments on IBMQ-Brisbane. The results show that CAZZO improves execution time by up to 37.44%, with an average improvement of 14.19%, across devices with 53 to 127 qubits. Real-device tests confirm CAZZO’s ability to reduce runtimes while preserving fidelity, outperforming pulse-based approaches.
Overall, our algorithm has achieved significant progress in optimizing quantum circuit execution efficiency and addressing the issue of ZZ crosstalk. Future work could focus on extending CAZZO’s capabilities to incorporate additional error mitigation techniques, such as dynamic decoupling, to further enhance fidelity while maintaining the efficiency gains. Additionally, exploring the potential for hybrid quantum-classical co-optimization could provide even more robust performance, particularly for circuits with specific noise characteristics or error models.
@@ Appendix A Detailed Data
This appendix provides partial detailed runtime comparisons between the CAZZO and baseline methods. The tables present execution times (in microseconds) across different benchmark datasets and hardware configurations.
@@
Benchmark ZZXSched ($\mu s$) CAZZO ($\mu s$) $\mathbf{\Delta}$ (%)
@@
Benchmark ZZXSched ($\mu s$) CAZZO ($\mu s$) $\mathbf{\Delta}$ (%)

### v2

@@ 7.2 RQ2 — Qubit Connectivity
Interestingly, our results suggest that CYCO performs particularly well in environments with Low Connectivity qubit architectures. The low connectivity structures, such as those found in IBMQ-Brisbane and Rigetti’s Aspen-M, tend to benefit more from intelligent gate scheduling, as the physical qubit interactions are constrained by the hardware topology. CYCO’s ability to optimize scheduling in such cases leads to better parallelism and resource utilization, as demonstrated by the greater speedups observed on these devices. In contrast, Linear Nearest Neighbor architectures [23], which inherently limit gate concurrency, show less dramatic improvements in execution time, though CYCO still provides notable gains.
@@ 7.3 RQ3 — Fidelity Maintenance
Figure 10 demonstrates CYCO’s ability to maintain computational fidelity while achieving significant speedups across five benchmark circuits on IBMQ Brisbane. For the bb84_n8 and qec_sm_n5 benchmarks, CYCO matches ZZXSched’s exact fidelity values. The lpn_n5 circuit reveals CYCO’s error resilience, maintaining 64% fidelity compared to the Standard scheduler’s 50%, a 14% relative improvement.
Our results disprove the assumption that reduced circuit duration necessarily increases error rates. CYCO maintains ZZXSched’s error suppression capability (calculated through fidelity ratio analysis) — achieving an optimal balance between speed and accuracy for practical quantum applications.
@@ 7.3.1 Trade-offs Between Density and Fidelity
Our analysis suggests a subtle trade-off between circuit density and fidelity. As CYCO increases the quantum circuit density by scheduling gates more compactly in time, it reduces the overall execution time, which benefits fidelity to the decreased exposure to decoherence. However, the denser packing of gates also increases the likelihood of crosstalk and other noise sources. The key takeaway from our experiments is that CYCO strikes a favorable balance, achieving significant reductions in execution time while maintaining high fidelity.
@@ 7.4 RQ4 — Scalability Analysis
Figure 11 presents a scalability comparison of CYCO versus ZZXSched across four different quantum architectures as the number of qubits increases. In smaller circuits, the gap between CYCO and ZZXSched is relatively narrow; however, as circuit size grows, CYCO demonstrates a consistently lower execution time. This trend indicates better scalability, suggesting that CYCO more effectively mitigates crosstalk and schedules gates in parallel without incurring significant overhead. Although the absolute times vary among architectures due to differences in gate durations and connectivity, CYCO maintains its advantage in each case.
@@ 9 Conclusions
We formally defined the cycle-aware ZZ crosstalk mitigation problem and proposed a corresponding optimization algorithm. CYCO balances ZZ crosstalk and parallelism while maximizing time resources in quantum circuits. It uses a new data structure, the Time and Distance Dependency Graph (TDDG), to capture dependencies and adjust gate execution based on varying durations. We assessed CYCO through simulations on 72 benchmarks from the QASMBench suite and real-device experiments on IBMQ-Brisbane. The results show that CYCO improves execution time by up to 37.44%, with an average improvement of 14.19%, across devices with 53 to 127 qubits. Real-device tests confirm CYCO’s ability to reduce runtimes while preserving fidelity, outperforming pulse-based approaches.
Overall, our algorithm has achieved significant progress in optimizing quantum circuit execution efficiency and addressing the issue of ZZ crosstalk. Future work could focus on extending CYCO’s capabilities to incorporate additional error mitigation techniques, such as dynamic decoupling, to further enhance fidelity while maintaining the efficiency gains. Additionally, exploring the potential for hybrid quantum-classical co-optimization could provide even more robust performance, particularly for circuits with specific noise characteristics or error models.
@@ Appendix A Detailed Data
This appendix provides partial detailed runtime comparisons between the CYCO and baseline methods. The tables present execution times (in microseconds) across different benchmark datasets and hardware configurations.
@@
Benchmark ZZXSched ($\mu s$) CYCO ($\mu s$) $\mathbf{\Delta}$ (%)
@@
Benchmark ZZXSched ($\mu s$) CYCO ($\mu s$) $\mathbf{\Delta}$ (%)